OpenStudy (anonymous):

I'm not sure I completely understand why the following differential / integral manipulation is mathematically correct.

OpenStudy (anonymous):

\[\int\limits_{0}^{t}L \frac{ di(t) }{ dt } i(t) dt = \int\limits_{0}^{i(t)}L i(t) di(t) \]

OpenStudy (anonymous):

What I don't understand is why the upper bound is i(t) and not t while the right and left hand side of the equation remain equivalent.

OpenStudy (anonymous):

This is from an electrical engineering book. Thus i(t) is the current as a function of time, thus t > 0

OpenStudy (anonymous):

I'm thinking it might be u-substitution

OpenStudy (anonymous):

\[u=i(t)\] \[du=\frac{ di(t) }{ dt }di(t)\] then change the limits